OpenStudy (anonymous):

Which ratio is equivalent to 8 : 32?

ganeshie8 (ganeshie8):

\(\large a : b\) is same \(\large \frac{a}{b}\)

OpenStudy (anonymous):

A. 1 : 16 B. 1 : 24 C. 1 : 4 D. 1 : 8

ganeshie8 (ganeshie8):

\(\large \color{red}{8 : 32} = \frac{8}{32} = \frac{8\times 1 }{8 \times 4 } = \frac{\not 8\times 1 }{\not8 \times 4 } = \frac{ 1 }{ 4 } = \color{red}{1 : 4}\)